how do i solve the system by graphing.
x – y = 5
x + y = 3

Answer by checkley71(8403)

SET X=0 AND SOLVE 0-Y=-5
SET Y=0 AND SOLVE X-0=-5
PLOT THESE TWO POINTS AND DRAW A LINE THROUGH THEM (-5,-5)
-----------------------------------------
SOLVE THE SECOND EQUATION THE SAME WAY
X+0=3
X=3
0+Y=3
Y=3 PLOT THESE TWO POINTS AND DRAWA LINE THROUGH THEM (3,3)
---------------------------------------------
WHERE THESE TWOLINE INTERSECT IA THE SOLUTION (4,-1)